Circular No. 6456

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1996 N2 (ELST-PIZARRO) On Aug. 7 Eric W. Elst, Royal Observatory, Uccle, reported his dis-

covery of a possible comet on mid-July exposures by Guido Pizarro with the 1.0-m Schmidt at the European Southern Observatory. Further ESO Schmidt plates were then obtained, and on Aug. 19, with the help of orbit computations by the undersigned, Elst was able to identify the object on them. The continued presence of a tail seemingly confirms the object as a “comet”, even though the orbit is completely that of a main-belt minor planet with the implied long-term stability.

1996 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 14.27639 22^{h}00^{m}05.^{s}53 −11^{o}16′51′′.5 18.3 Pizarro
14.30278 22 00 04.85 −11 16 55.5  _{”}
16.29514 21 59 21.38 −11 21 23.5  _{”}
16.32153 21 59 20.61 −11 21 28.3  _{”}
Aug. 8.27986 21 45 48.76 −12 42 18.0  _{”}
8.30625 21 45 47.50 −12 42 24.7  _{”}
8.97986 21 45 18.07 −12 45 23.6  Pollas
8.99028 21 45 17.57 −12 45 25.3  _{”}
9.23889 21 45 06.50 −12 46 25.9 18.0 Pizarro
9.26528 21 45 05.20 −12 46 32.7  _{”}
17.95139 21 38 33.54 −13 24 40.5  Pollas
17.97222 21 38 32.60 −13 24 43.2  _{”}
20.0271 21 37 00.3 −13 33 39  Elst
20.89057 21 36 22.30 −13 37 19.7 17.1 Tichá
21.03161 21 36 15.85 −13 37 55.6  _{”}
21.26023 21 36 05.96 −13 38 53.1 18.2 Offutt
21.28061 21 36 04.98 −13 38 58.7 18.5  _{”}

G. and O. Pizarro (European Southern Observatory). 1.0-m Schmidt. Mea-

surer E. W. Elst. Narrow tail > 1′ long in southwestern quadrant. C. Pollas (Caussols). 0.9-m Schmidt. Measurer E. W. Elst. E. W. Elst (Uccle). 0.8-m Schmidt + CCD. Tail in p.a. 250^{o}. J. Tichá, M. Tichý and Z. Moravec (Kleť). 0.57-m f/5.2 reflector + CCD.

Tail > 3′ long to west-southwest. No evident coma. W. Offutt (Cloudcroft). 0.6-m f/7 Ritchey-Chrétien + CCD. Bright, sharp-

ly-defined, structureless tail > 3′.9 long (off edge of field) in p.a. 251^{o}. Head completely stellar, no discernible coma.

1996 August 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden